# supplementary

### v2

@@ Abstract
Keywords: Parallel tempering; Mode-jumping proposals; Curved exponential family; Seemingly-unrelated regression; Panel data; Raman spectroscopy
@@ 1 Introduction
The standard approach to sampling from a multivariate Bayesian posterior density is Markov Chain Monte Carlo (MCMC). There is a wealth of established literature and methodology enabling the construction of a Markov chain with prescribed invariant distribution (e.g., Robert and Casella, 2013). Arguably the easiest way to construct such a chain is by using a localised proposal mechanism e.g., random walk Metropolis (Metropolis et al., 1953), Metropolis-adjusted Langevin (Rossky et al., 1978; Roberts and Tweedie, 1996), or hybrid methods formed from concatenating local moves such as Hamiltonian Monte Carlo (Duane et al., 1987). Such proposals can then be accepted or rejected according to the Metropolis-Hastings acceptance probability (Hastings, 1970), ensuring reversibility of the chain.
When the target exhibits multimodality, the majority of MCMC algorithms that use localised proposals will fail to explore the state space, leading to biased estimation of posterior expectations from the sample output (Roberts et al., 2001; Mangoubi et al., 2018). Indeed, when the target distribution is high-dimensional or the posterior arises from a complex Bayesian modeling context then it becomes very hard to know whether the posterior is multimodal. Therefore it is important to develop methods that are robust but also scale well to high-dimensional settings.
The most popular MCMC methods designed to explore multimodal distributions have traditionally been based on tempering approaches and building on the seminal contributions of Geyer (1991); Marinari and Parisi (1992); Geyer and Thompson (1995). However, despite the success of these methods, tempering approaches often fail in high-dimensional settings, and where individual modes have different variances and/or display asymmetries (Woodard et al., 2009a, b).
This paper introduces a new algorithm, the Annealed Leap Point Sampler (ALPS), that overcomes many of these disadvantages. It is designed to sample from a class of smooth densities on $\mathbb{R}^{d}$ that may exhibit multimodality. The ALPS construction uses an annealing approach, where contrary to traditional approaches it relies on raising the density to powers greater than one. It also adapts ideas from established methods in the literature designed to mitigate the worst effects of multimodality in multi-dimensional settings, such as mode-jumping proposals (Tjelmeland and Hegstad, 2001) and weight-stabilised tempering (Tawn and Roberts, 2018; Tawn et al., 2020) that incorporate non-local information about the target.
@@ 2 Background and Motivation
However the most successful and widely adopted approach to the problem are based on state-space augmentation techniques. These stem from two seminal, closely related but independently discovered algorithms: the parallel tempering (PT) algorithm (Geyer, 1991) and its closely-related cousin, the simulated tempering algorithm (Marinari and Parisi, 1992). The principle behind these methods is that a Markov chain can achieve global exploration when targeting a tempered version of the density ($\pi^{\beta}$ for some $10<\beta<$ 1). Therefore the strategy is to construct an MCMC algorithm on an augmented space, ${\mathbb{R}^{d}\times\Delta}$ where $\Delta=\{\beta_{j},j=0,\ldots n\}\subset(0,1]$, with $\beta_{0}=1$ denotes the temperature of interest. Since $\pi$ is assumed to be multi-modal, direct Markov chain transitions between its modes are likely to be very difficult. However, indirect paths between these modes can be achieved by moving to hotter temperatures (smaller $\beta$) where inter-mode transitions are feasible, and then moving temperature back to $\beta_{0}=1$.
@@
We shall show that this ambitious plan can be accomplished by giving detailed methodology and supporting theory that demonstrates its robustness in high dimensional contexts. In particular, we shall see that this annealing approach can be carried out without an increase in computational complexity in comparison to standard tempering methods.
@@ 3 Components of ALPS
Crucial to the design of the algorithm will be the following Gaussian mixture distribution.
[…]
@@
If the target distribution has modes that locally resemble Gaussian distributions, it can naturally be approximated effectively by a Gaussian mixture like the one in (2), where $m$ would correspond to the number of local maxima. In such cases, provided the Gaussian mixture parameters can be derived, the problem simplifies considerably. Indeed, by combining localized MCMC moves with independence sampler moves drawn from the proposal distribution given in (2), one can design a Markov chain that would have very rapid convergence.
@@ 3.1 Hessian-Adjusted Tempered (HAT) Distributions
Woodard et al. (2009a) shows that when modes of a multimodal target exhibit heterogeneity, particularly with regards to their scales, then these power-based target distributions are a poor choice as they lead to torpid mixing of PT when the dimensionality of the problem grows. Unless the distributions are altered, annealing is subject to similar challenges of torpid mixing. The chains targeting the most annealed distributions would become increasingly confined to the dominant mode containing the global maximum as in simulated annealing (Aarts and Korst, 1988).
@@
First, we introduce an approximate partition of the state space into domains (or basins) of attraction, in the sense of Tjelmeland and Hegstad (2001); Zhou (2011a). In this context, the domain of attraction $\mathcal{D}_{j}\subseteq\mathcal{X}$ associated with mode $\mu_{j}$ is defined as the set of points from which the gradient flow dynamics (i.e., the solutions to the gradient flow ODE) converge to $\mu_{j}$, as introduced in Zhou (2011a). Given the Gaussian mixture distribution $\pi_{M,S,W}$, we derive an approximation of these domains, based on a modified Mahalanobis distance criterion (dependent on $M,S,W$, although we omit this explicit dependence for notational convenience).
@@
This assignment yields a partitioning that matches the gradient flow definition of domains of attraction in the specific case that the target distribution is itself a Gaussian mixture. Via annealing the modes approach local Gaussianity and so the precision of our approximation to these domains becomes increasingly accurate. Using this partition, we correspondingly define the Hessian Adjusted Tempered (HAT) distributions that will be used as our auxiliary densities $\pi_{\beta_{j}}$.
@@
Theoretical justification for these distributions can be found in the supplementary material in Appendix LABEL:sec:Weightstabilised. In short this form of the target distribution can mitigate some of the most undesirable weight stabilisation issues exhibited by traditional tempering approaches.
@@ 3.2 Choice of inverse temperatures
Appealing to the theoretical contributions of Tawn and Roberts (2018), Tawn and Roberts (2019) and Tawn et al. (2020) it is approximately optimal to select the spacings consecutively with $\beta_{j+1}=\beta_{j}+\epsilon_{j}$ where $\epsilon_{j}=\mathcal{O}\left(d^{-1/2}\right)$. Crucially, the optimal choice for the spacings corresponds to temperature swap acceptance rates between consecutive temperatures of 0.234. A practitioner can use this to tune the setup of the algorithm; even pursuing an adaptive strategy similar to Miasojedow et al. (2013).
@@ 3.3 Design of temperature swap moves
Within ALPS-PA we shall make extensive use of the QuanTA transformation (as originally introduced in Tawn and Roberts (2019)).
@@
Given a tempering schedule $\Delta$ with $\beta_{0}<\ldots<\beta_{n}$ and lists $M,S,W$, then define $T^{+}_{M,S,W}:\mathbb{R}^{d}\times\{\beta_{0},\ldots,\beta_{n-1}\}\to\mathbb{R}^{d}\times\{\beta_{1},\ldots,\beta_{n}\}$ by
$T^{+}_{M,S,W}(x,\beta_{i})=\left(\left(\frac{\beta_{i}}{\beta_{i+1}}\right)^{1/2}(x-\mu_{A_{x,\beta_{i}}})+\mu_{A_{x,\beta_{i}}},\ \beta_{i+1}\right)\ .$ (5)
Note that $T^{+}_{M,S,W}$ is injective (though not surjective) and so its inverse is well-defined with its domain given by the image of $T^{+}_{M,S,W}$.
@@
    $y_{i+1}=T^{+}_{M,S,W}(x_{i}^{t},\beta_{i})~{}~{}~{}\mbox{and}~{}~{}~{}y_{i}=T^{+}_{M,S,W}(x_{i+1}^{t},\beta_{i+1})$
@@
    Check for reversibility of the transformation (as in rare cases the transformation may have crossed a boundary voiding the reversibility immediately). Immediately reject the proposed swap if
@@
    If (c) does not lead to rejection then compute
    $\alpha=\min\left\{1,\frac{\pi_{\beta_{i}}(y_{i})\pi_{\beta_{i+1}}(y_{i+1})}{\pi_{\beta_{i}}(x_{i})\pi_{\beta_{i+1}}(x_{i+1})}\right\}.$
@@
The transformation is more effective when the modes exhibit a more Gaussian structure. For robustness, it is wise to mix in traditional swap moves around the target temperature particularly in cases when the modes are highly non-Gaussian.
@@ 3.4 Choice of within-mode transitions
To increase robustness, we recommend employing sampling schemes that explicitly exploit local geometric properties which can differ significantly between modes, e.g. pre-conditioned Random Walk Metropolis. There are many established results for tuning the within-temperature marginal updates to the chain performed in PA component, e.g. Roberts et al. (2001).
@@
The collections $M,S$ and $W$ will be estimated via Laplace approximations as detailed in Section 3.6. If sufficient annealing has taken place so that the Laplace approximations for the local modes are reasonably accurate then the weight approximations will closely match the regional masses in the $\beta_{\text{max}}$-level HAT target. In that case, the independence sampler constructed above should be able to achieve high acceptance rates.
Of course, any finite amount of annealing leaves some modal asymmetry and thus prevents acceptance rates being equal to 1. In general, one must anneal more as the dimensionality of the problem increases before the Laplace approximation becomes accurate. The question of how large $\beta_{\text{max}}$ should be with respect to the dimensionality of the problem to achieve sufficiently good performance is addressed in Section 5.
@@ 3.6 Design of mode-finding procedure
ALPS-EC constructs a Markov chain that has an invariant distribution given by $[\pi(\cdot)]^{\beta_{\text{hot}}}$. If the inverse-temperature level, $\beta_{\text{hot}}$, is sufficiently low then, as in the standard PT algorithm, the Markov chain can explore the entire state space rapidly. After a pre-specified number of steps of the Markov chain, one performs a local optimisation (this work uses Quasi-Newton optimisation) initialised from the current location of the chain. If the optimum attained is a new mode point not previously encountered, one adds it to the collection of discovered modes.
@@
We do not claim that our approach to mode finding is optimal, nor do we provide theoretical guarantees. The process of identifying modes is computationally intensive and highly problem-dependent; no single approach can be expected to work best in all scenarios. Improving the performance of the mode-finding procedure is a focus of further work beyond the scope of this paper. In practice, empirical evidence suggests that ALPS-EC should be adapted to the specific characteristics of each problem.
@@ 4 ALPS : The procedure
      For $10\leq j\leq n-$ 1, draw $x_{j}^{t+1}\sim P_{\beta_{j}}(x_{j}^{t},.)$ for local exploration, see (3.1) and (3.4);
@@
    Store the subsequent sample $X^{t+1}=(x_{0}^{t+1},\ldots,x_{n}^{t+1})$, go to step (a).
@@
A comprehensive pseudocode version of ALPS, suitable for direct implementation, is presented in Appendix LABEL:sec:PseudoCode of the online supplementary material.
@@ 5 Theoretical Underpinnings of ALPS
Choosing $\beta_{\text{max}}$ extremely large to overcome this has a few issues. On a practical level, it may result in computational instability as modes collapse towards point masses. The HAT targets are only approximately preserving mass so their performance may become degenerate for excessively large $\beta$. Finally, there is extra computational burden as one would need more (unnecessary) intermediate temperature levels in the PA component. Hence, one would like to choose $\beta_{\text{max}}$ sufficiently large to achieve non-degenerate acceptance rates but not so large that it makes round trips infrequent.

### v3

# Annealed Leap-Point Sampler for Multimodal Target Distributions

@@ 1 Introduction
The standard approach to sampling from a multivariate Bayesian posterior density is Markov Chain Monte Carlo (MCMC). There is a wealth of established literature and methodology enabling the construction of a Markov chain with prescribed invariant distribution (e.g., Robert and Casella, 2004). Arguably the easiest way to construct such a chain is by using a localised proposal mechanism e.g., random walk Metropolis (Metropolis et al., 1953), Metropolis-adjusted Langevin (Rossky et al., 1978; Roberts and Tweedie, 1996), or hybrid methods formed from concatenating local moves such as Hamiltonian Monte Carlo (Duane et al., 1987). Such proposals can then be accepted or rejected according to the Metropolis-Hastings acceptance probability (Hastings, 1970), ensuring reversibility of the chain.
When the target exhibits multimodality, the majority of MCMC algorithms that use localised proposals will fail to explore the state space, leading to biased estimation of posterior expectations from the sample output (Roberts and Rosenthal, 2001; Mangoubi et al., 2018). Indeed, when the target distribution is high-dimensional or the posterior arises from a complex Bayesian modeling context then it becomes very hard to know whether the posterior is multimodal. Therefore it is important to develop methods that are robust but also scale well to high-dimensional settings.
The most popular MCMC methods designed to explore multimodal distributions have traditionally been based on tempering approaches and building on the seminal contributions of Geyer (1991); Marinari and Parisi (1992); Geyer and Thompson (1995). However, despite the success of these methods, tempering approaches often fail in high-dimensional settings, and where individual modes have different variances and/or display asymmetries (Woodard et al., 2009b).
This paper introduces a new algorithm, the Annealed Leap-Point Sampler (ALPS), that overcomes many of these disadvantages. It is designed to sample from a class of smooth densities on $\mathbb{R}^{d}$ that may exhibit multimodality. The ALPS construction uses an annealing approach, where contrary to traditional approaches it relies on raising the density to powers greater than one. It also adapts ideas from established methods in the literature designed to mitigate the worst effects of multimodality in multi-dimensional settings, such as mode-jumping proposals (Tjelmeland and Hegstad, 2001) and weight-stabilised tempering (Tawn and Roberts, 2018; Tawn et al., 2020) that incorporate non-local information about the target.
@@ 2 Background and Motivation
However the most successful and widely adopted approach to the problem are based on state-space augmentation techniques. These stem from two seminal, closely related but independently discovered algorithms: the parallel tempering (PT) algorithm (Geyer, 1991) and its closely-related cousin, the simulated tempering algorithm (Marinari and Parisi, 1992). The principle behind these methods is that a Markov chain can achieve global exploration when targeting a tempered version of the density ($\pi^{\beta}$ for some $0<\beta<1$). Therefore the strategy is to construct an MCMC algorithm on an augmented space, $\mathbb{R}^{d}\times\Delta$} where $\Delta=\{\beta_{j},j=0,\ldots n\}\subset(0,1]$, with $\beta_{0}=1$ denoting the temperature of interest. Since $\pi$ is assumed to be multi-modal, direct Markov chain transitions between its modes are likely to be very difficult. However, indirect paths between these modes can be achieved by moving to hotter temperatures (smaller $\beta$) where inter-mode transitions are feasible, and then moving temperature back to $\beta_{0}=1$.
@@
We shall show that this ambitious plan can be accomplished by giving detailed methodology and supporting theory that demonstrates its robustness in high dimensional contexts. In particular, we demonstrate that the proposed annealing strategy achieves linear mixing-time scaling in dimension, addressing regimes in which standard tempering methods may exhibit exponentially slow convergence, e.g. Woodard et al. (2009b).
This linear behaviour persists even for well-separated, skewed modes and heterogeneous covariance structures, ensuring efficient mixing as dimension increases. The resulting gain in Markov chain convergence shifts the complexity from exponential to polynomial, at the expense of a modest increase in per-iteration computational cost, preserving polynomial-time scalability.
@@ 3 Components of ALPS
A central component in the design of the algorithm is the following Gaussian mixture distribution. Given a list of vectors $M=\{\mu_{j},\ 1\leq j\leq m\}$, a list of symmetric positive-definite matrices $S=\{\Sigma_{j},\ 1\leq j\leq m\}$, and a list of positive weights summing to unity $W=\{w_{1},\ldots,w_{m}\}$, we set
@@
If the target distribution has modes that locally resemble Gaussian distributions, it can naturally be approximated effectively by a Gaussian mixture like the one in (2), where $m$ would correspond to the number of local maxima. In such cases, provided the Gaussian mixture parameters can be derived, the problem simplifies considerably. Indeed, by combining localized MCMC moves with independence sampler moves drawn from the proposal distribution given in (2), one can design a Markov chain that would yield very rapid convergence.
@@ 3.1 Hessian-Adjusted Tempered (HAT) Distributions
Woodard et al. (2009b) shows that when modes of a multimodal target exhibit heterogeneity, particularly with regards to their scales, then these power-based target distributions are a poor choice as they lead to torpid mixing of PT when the dimensionality of the problem grows. Unless the distributions are altered, annealing is subject to similar challenges of torpid mixing. The chains targeting the most annealed distributions become increasingly restricted to the dominant modes corresponding to global maxima, as in simulated annealing (Aarts and Korst, 1988).
@@
First, we introduce an approximate partition of the state space into domains (or basins) of attraction, in the sense of Tjelmeland and Hegstad (2001); Zhou (2011a). In this context, the domain of attraction $\mathcal{D}_{j}\subseteq\mathcal{X}$ associated with mode $\mu_{j}$ is defined as the set of points from which the gradient flow dynamics (i.e., the solutions to the gradient-flow ODE) converge to $\mu_{j}$, as introduced in Zhou (2011a). Given the Gaussian mixture distribution $\pi_{M,S,W}$, we derive an approximation of these domains, based on a modified Mahalanobis distance criterion (dependent on $M,S,W$, although we omit this explicit dependence for notational convenience).
@@
This assignment produces a partition consistent with the gradient-flow definition of domains of attraction when the target distribution is a Gaussian mixture. Through annealing, the modes approach local Gaussianity, and the approximation of these domains becomes increasingly accurate.. Using this partition, we follow Tawn et al. (2020, Def. 2) and restate the Hessian Adjusted Tempered (HAT) distributions, which we adopt as our auxiliary densities $\pi_{\beta_{j}}$.
@@
Theoretical justification for these distributions can be found in the supplementary material in Appendix B.1. In short, this form of the target distribution can mitigate some of the most undesirable weight stabilisation issues exhibited by traditional tempering approaches.
@@ 3.2 Choice of inverse temperatures
Appealing to the theoretical contributions of Tawn and Roberts (2018), Tawn and Roberts (2019) and Tawn et al. (2020), it is approximately optimal to select the spacings consecutively with $\beta_{j+1}=\beta_{j}+\epsilon_{j}$ where $\epsilon_{j}=\mathcal{O}\left(d^{-1/2}\right)$. Crucially, the optimal choice for the spacings corresponds to temperature swap acceptance rates between consecutive temperatures of 0.234. A practitioner can use this to tune the setup of the algorithm; even pursuing an adaptive strategy similar to Miasojedow et al. (2013).
@@ 3.3 Design of temperature swap moves
Within ALPS-PA we shall make extensive use of the QuanTA transformation (as originally introduced in Tawn and Roberts, 2019).
@@
Given a tempering schedule $\Delta$ with $\beta_{0}<\ldots<\beta_{n}$ and lists $M,S,W$, then define $T^{+}_{M,S,W}:\mathbb{R}^{d}\times\{\beta_{0},\ldots,\beta_{n-1}\}\to\mathbb{R}^{d}\times\{\beta_{1},\ldots,\beta_{n}\}$ and $T^{-}_{M,S,W}:\mathbb{R}^{d}\times\{\beta_{1},\ldots,\beta_{n}\}\to\mathbb{R}^{d}\times\{\beta_{0},\ldots,\beta_{n-1}\}$ by
$T^{+}_{M,S,W}(x,\beta_{i})$ $=\left(\left(\frac{\beta_{i}}{\beta_{i+1}}\right)^{1/2}(x-\mu_{A_{x,\beta_{i}}})+\mu_{A_{x,\beta_{i}}},\ \beta_{i+1}\right),$ (5a)
$T^{-}_{M,S,W}(x,\beta_{i})$ $=\left(\left(\frac{\beta_{i}}{\beta_{i-1}}\right)^{1/2}(x-\mu_{A_{x,\beta_{i}}})+\mu_{A_{x,\beta_{i}}},\ \beta_{i-1}\right).$ (5b)
Note that $T^{+}_{M,S,W}$ and $T^{-}_{M,S,W}$ are both injective (though not surjective) and so their inverse are well-defined within the domains given by their respective images.
@@
    $y_{i+1}=\left(T^{+}_{M,S,W}(x_{i}^{t},\beta_{i})\right)_{1}~~~\mbox{and}~~~y_{i}=\left(T^{-}_{M,S,W}(x_{i+1}^{t},\beta_{i+1})\right)_{1}$
@@
    Check for reversibility of the transformation (as in rare cases it may have crossed a boundary, immediately voiding reversibility). Reject the proposed swap if
@@
    If 3. does not lead to rejection then compute
    $\alpha=\min\left\{1,\frac{\pi_{\beta_{i}}(y_{i})\pi_{\beta_{i+1}}(y_{i+1})}{\pi_{\beta_{i}}(x_{i}^{t})\pi_{\beta_{i+1}}(x_{i+1}^{t})}\right\}.$
@@
The transformation is more effective when the modes exhibit approximately Gaussian structure. For robustness in highly degenerate cases where QuanTA swaps may start to fail or, for instance, when modes are highly non-Gaussian, occasional traditional swap moves may optionally be included near the target temperature. In practice, we did not observe such behaviour, and empirical evidence suggests that this safeguard is not necessary.
@@ 3.4 Choice of within-mode transitions
To increase robustness, we recommend employing sampling schemes that explicitly exploit local geometric properties which can differ significantly between modes, e.g. pre-conditioned Random Walk Metropolis. There are many established results for tuning the within-temperature marginal updates to the chain performed in PA component, e.g. Roberts and Rosenthal (2001).
@@
The collections $M$, $S$ and $W$ will be estimated via Laplace approximations as detailed in Section 3.6. If adequate annealing has taken place so that the Laplace approximations for the local modes are reasonably accurate then the weight approximations will closely match the regional masses in the $\beta_{\text{max}}$-level HAT target. In that case, the independence sampler constructed above should be able to achieve high acceptance rates.
Of course, any finite amount of annealing leaves some modal asymmetry and thus prevents acceptance rates being equal to 1. In general, one must anneal more as the dimensionality of the problem increases before the Laplace approximation becomes accurate. The question of how large $\beta_{\text{max}}$ should be with respect to the dimensionality of the problem to achieve sufficiently good performance is addressed in Section 5. In practice, to choose such a value of $\beta_{\text{max}}$, any target acceptance rate $a\in(0,1)$ at the cold level achieves the corresponding complexity guarantees described in Section 6. A natural choice is therefore $a=0.5$, which can be increased if the resulting hopping rate is not sufficiently high for the problem at hand. Our empirical results do not indicate a clear preference for a particular value of $a$; indeed, Section 7 presents experiments covering acceptance rates ranging from 0.25 to 0.9. Nevertheless, for robustness, a practitioner may choose a higher target acceptance rate, such as 0.9, to ensure frequent hopping. This does not incur computational burden under the chosen temperature schedule in Section 3.3.
@@ 3.6 Design of mode-finding procedure
ALPS-EC constructs a Markov chain that has an invariant distribution given by $\pi(\cdot)]^{\beta_{\text{hot}}}$. If the inverse-temperature level, $\beta_{\text{hot}}$, is sufficiently low then, as in the standard PT algorithm, the Markov chain can explore the entire state space rapidly. After a pre-specified number of steps of the Markov chain, one performs a local optimisation (this work uses Quasi-Newton optimisation) initialised from the current location of the chain. If the optimum attained is a new mode point not previously encountered, one adds it to the collection of discovered modes.
@@
We do not claim that our approach to mode finding is optimal, nor do we provide theoretical guarantees. The process of identifying modes is computationally intensive and highly problem-dependent; no single approach can be expected to work best in all scenarios. Improving the performance of the mode-finding procedure is a focus of further work beyond the scope of this paper. In practice, it may be beneficial to adapt ALPS–EC to the characteristics of the problem at hand. For instance, in machine-learning settings, where objective functions are often evaluated using minibatches and are therefore noisy, stochastic gradient-based methods such as SGD or Adam (Kingma and Ba, 2017) are natural candidates for identifying regions of high density.
@@ 4 ALPS : The procedure
      For $0\leq j\leq n-1$, draw $x_{j}^{t+1}\sim P_{\beta_{j}}(x_{j}^{t},.)$ for local exploration, see (3.1) and (3.4);
@@
    Store the subsequent sample $X^{t+1}=(x_{0}^{t+1},\ldots,x_{n}^{t+1})$, go to step 1.
@@
A comprehensive pseudocode version of ALPS, suitable for direct implementation, is presented in Appendix C of the online supplementary material.
@@ 5 Theoretical Underpinnings of ALPS
Choosing $\beta_{\text{max}}$ extremely large to overcome this has a few issues. On a practical level, it may result in computational instability as modes collapse towards point masses. Furthermore, the HAT targets provide only approximate mass preservation, and care is therefore required when considering extremely large values of the inverse temperature; see Tawn et al. (2020). Finally, there is extra computational burden as one would need more (unnecessary) intermediate temperature levels in the PA component. Hence, one would like to choose $\beta_{\text{max}}$ sufficiently large to achieve non-degenerate acceptance rates but not so large that it makes round trips infrequent.